If a swimming pool has a perimeter of 18 meters and covers an area of 18 square meters what are its dimensions?
Answer by ankor@dixie-net.com(22740)

If a swimming pool has a perimeter of 18 meters and covers an area of 18 square meters what are its dimensions?
:
Perimeter
2L + 2W = 18
simplify, divide by 2
L + W = 9
L = (9-W)
:
Area
L * W = 18
replace L with (9-W)
W(9-W) = 18
A quadratic equation
-W^2 + 9W - 18 = 0
Easier to factor if coefficient of W^2 is positive, mult eq by -1
W^2 - 9W + 18 = 0
Factors to
(W-3)(W-6) = 0
Two solutions
W = 3, then L = 6
W = 6, then L = 3